OpenStudy (anonymous):

Which ordered pair is a solution to the equation 3x + 5y = 9? A. (–2, –3) B. (–2, 3) C. (2, –3) D. (2, 3)

OpenStudy (anonymous):

I believe its D

OpenStudy (anonymous):

how did you get the answer? I like to know because i need to know how to work it out for myself when the test comes

OpenStudy (texaschic101):

sub in your answer choices and see which one satisfies the equation. 3x + 5y = 9 (-2,3) x = -2 and y = 3 now sub 3(-2) + 5(3) = 9 -6 + 15 = 9 9 = 9 (correct) answer is B

OpenStudy (anonymous):

ok thanks so much!

OpenStudy (texaschic101):

do you understand how I did that ?

OpenStudy (anonymous):

OOOOH. my bad. /.o

OpenStudy (texaschic101):

each of your points is set up (x,y). Just sub in the number for x and the number for y in the equation and find which one is correct